OpenStudy (anonymous):

AWARD AND MEDALLLLL Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (debbieg):

You need some variables here. Let D=# of dimes, and Q=# of quarters. Can you write an equation to model the fact that the total number of coins is 46?

OpenStudy (anonymous):

D+Q=46

OpenStudy (debbieg):

Good! Now we have 2 variables, so we need 2 equations. How about an equation that models the fact that the total value of the coins is 7.00?

OpenStudy (anonymous):

D(.10)+Q(.25)=7.00

OpenStudy (debbieg):

Excellent! Now, you probably don't like those pesky decimals in that equation any more than I do, do you? I would probably eliminate the decimals in that 2nd equation. How can we do that?

OpenStudy (debbieg):

And what do you get when you do?

OpenStudy (anonymous):

uhm, multiply each number by 100. D(10)+Q(25)=700

OpenStudy (debbieg):

Look at you go! You hardly need my help. :) Yes, what that did is basically converted that 2nd equation from being in terms of dollars to being in terms of cents. Now you have these 2 equations. D+Q=46 10D+25Q=700 2 equations.... 2 unknowns.... do you know what to do from here? That first one is conveniently easy to solve for either variable, and then plug into the 2nd equation.

OpenStudy (anonymous):

ooops nevrmind

OpenStudy (debbieg):

hmmm.... I don't think so, how did you get to that point? Did you solve the first equation for one of the variables? (I suggest solving for D in terms of Q, but either way will work)

OpenStudy (anonymous):

I'm kind of stuck :| Sorry

OpenStudy (debbieg):

That's ok, let's take it apart a little bit now. You have these 2 equations: D+Q=46 10D+25Q=700 Let's solve that first one for D. In other words, shuffle the equation around so you have D=?

OpenStudy (anonymous):

D=46-Q

OpenStudy (debbieg):

Good. That's "D in terms of Q" because it expresses D based on what Q is. Now you can take THAT expression for D, and plug it into the OTHER equation where you see D. 10(D)+25Q=700 Replace the D with (46-Q)

OpenStudy (debbieg):

That will give you an equation in ONE variable. Just simplify the expressions and isolate the variable Q, and you will get a value for Q. Once you have that value for Q, you can plug THAT back into your equation that gives D in terms of Q, and you will get D. And then of course, you can CHECK your answer, making sure it fits what the original problem said. :)

OpenStudy (anonymous):

10(46-Q)+25Q=700 460-10Q+25Q=700 460-35Q=700 700-460=240 35Q=240 I come up with a weird/ long decimal....What did I do wrong?

OpenStudy (debbieg):

What is -10Q+25Q? :)

OpenStudy (anonymous):

ooh I forgot that it was negative

OpenStudy (debbieg):

That will make a big difference. :)

OpenStudy (anonymous):

10(46-Q)+25Q=700 460-10Q+25Q=700 460-15Q=700 700-460=240 240/15= 16 Q=16

OpenStudy (debbieg):

Great! Now you know Q, so how do you get D?

OpenStudy (anonymous):

Subtract 240 from 700 and divide it by 10

OpenStudy (anonymous):

Wait thats not right

OpenStudy (debbieg):

haha... nooo.... you are making it too hard. You KNOW Q. You want D. Do you have an equation at your fingertips that... oh, I don't know.... give you D in terms of what Q is? Hm??

OpenStudy (debbieg):

(up approximately 11 posts) :)

OpenStudy (debbieg):

You ok? Figure it out?

OpenStudy (anonymous):

Oops sorry

OpenStudy (anonymous):

there are 16 quarters and 10 dimes

OpenStudy (debbieg):

How did you get 10 for the dimes? And does that add up to the amount of money that there is supposed to be?

OpenStudy (debbieg):

You have an equation for dimes, given quarters: D-46-Q So since you know quarters, that equation will let you figure the dimes. Then the total coins should add to ----- how many? And the total value of the coins should be ---- how much? Neither of these ^^^ checks will work with 16 quarters and 10 dimes.